There's a large broadband increase in the MC_F spectrum. I'm not totally sure it's real--it could be some weird bit-swapping thing. I've tried soft reboots of c1susvme2 and c1iovme, which haven't helped. In any case, it seems like this is preventing any locking success today. Last night it was fine.
Rebooting c1iovme (by keying off the crate, waiting 30 seconds, and then keying it back on and restarting) has resolved this. The frequency noise is back to the 'usual' trace.